# Does Your Multi-modal LLM Truly See the Diagrams in Visual Math Problems?

### v1

@@ 1 Introduction
With the substantial advances of big data and computational power, Large Language Models (LLMs) brown2020language ; albert24mixtral ; touvron2023llama ; touvron2023llama2 ; vicuna2023 , such as ChatGPT OpenAI2023ChatGPT and GPT-4 OpenAI2023GPT4TR , have emerged as a central point of interest in both industry and academia. To broaden their applicability across diverse contexts, Multi-modal Large Language Models (MLLMs) zhang2024llamaadapter ; gao2023llamaadapterv2 ; su2023pandagpt ; chen2023minigpt ; ye2023mplugowl have recently become a fast-evolving track, exemplified by the latest GPT-4V openai2023gpt4v , Gemini team2023gemini , and the open-source LLaVA liu2023llava ; liu2023improvedllava ; liu2024llavanext and SPHINX lin2023sphinx ; gao2024sphinx . Concurrently, a diverse array of evaluation benchmarks fu2023mme ; liu2023mmbench ; Li2023SEEDBenchBM ; fu2023challenger ; sun2024journeydb are curated to assess their visual comprehension performance across different domains. Notably, the capability to solve mathematical problems involving diagrams serves as a critical measure, offering insights into the multi-modal logical thinking prowess of MLLMs. This task demands MLLMs to accurately decode the visual elements within input diagrams (characters and figures), and correlate them with the condition specified by textual questions for mathematical reasoning. Previous efforts lu2021inter ; seo2015solving , e.g., GeoQA chen2021geoqa ; cao2022augmented and UniGeo chen2022unigeo , concentrate on the challenging geometric problems, while the recent MathVista Lu2023MathVistaEM and MMMU yue2023mmmu expand the scope to encompass broader disciplines, including functions, charts, and scientific problems.
Figure 1: (a) We showcase three examples of Text Redundancy (highlighted in red) within existing visual math benchmarks chen2021geoqa ; Lu2023MathVistaEM ; yue2023mmmu . (b) We report an ablation study by respectively removing the redundant texts and input diagrams on 120 randomly selected problems, for closed-sourced openai2023gpt4v ; team2023gemini ; bai2023qwen and open-sourced gao2024sphinx ; liu2024llavanext ; dong2024internlm MLLMs.
@@
Figure 2: Comparison of Visual Mathematical Reasoning by Three MLLMs. Despite the incorrect final answer, GPT-4V openai2023gpt4v , Gemini-Pro team2023gemini , and SPHINX-MoE gao2024sphinx exhibit different levels of quality in the intermediate reasoning process.
@@
Second, to rigorously assess the visual Chain-of-Thought (CoT) capabilities wei2022chain , we propose a CoT Evaluation strategy for the step-by-step reasoning assessment of MLLMs. For each model’s output, we leverage GPT-4 to first extract several crucial steps exclusively from the solving process, deliberately omitting the input of the question and answer. This approach aims to mitigate the bias towards GPT-4’s inherent question-answering propensities. Then, the corresponding question, diagram, and ground-truth answer are fed into GPT-4 to evaluate each identified critical step, and provide detailed error analysis. Finally, the overall score is obtained by considering every single step within reasoning. Note that, we do not pre-define a ground-truth key-step template, since each math problem may encompass a variety of solution pathways, and different MLLMs tend to exhibit variable reasoning lengths. With CoT scoring, MathVerse showcases a fine-grained evaluation of the intermediate logical deduction of MLLMs, demonstrating visual mathematical CoT capabilities.
@@
We conduct extensive experiments on MathVerse with popular closed-source openai2023gpt4v ; bai2023qwen ; team2023gemini and open-source liu2023improvedllava ; liu2024llavanext ; dong2024internlm ; gao2024sphinx MLLMs. Comparing different problem versions, we unveil that, most existing MLLMs struggle to understand math diagrams, relying heavily on textual questions. Therein, GPT-4V openai2023gpt4v achieves the best overall performance across different problem versions and subjects. Surprisingly, some of the MLLMs even attain much higher results without the diagram input, e.g., +5.1% for Qwen-VL-Max bai2023qwen and +5.6% for InternLM-XComposer2 dong2024internlm . With the fine-grained error analysis produced by our CoT evaluation strategy, we demonstrate such results are due to their deficient visual encoding capacity for mathematical diagrams, which instead acts as a distraction for problem-solving. In contrast, GPT-4V and ShareGPT4V Chen2023ShareGPT4VIL demonstrate relatively better comprehension of the visual content for mathematical reasoning. Our experimental results suggest that inadequate mathematical visual interpretation capabilities represent the most significant impediment for MLLMs in addressing multi-modal math problems, indicating substantial potential for advancement.
@@ 2 MathVerse
[…]
@@ 2.1 Visual Math Dataset
MathVerse comprises a total of 2,612 visual math problems, which contribute to the final created 15K test samples. Detailed statistics for data composition are presented in Table 4. This meticulously collected dataset covers three fundamental math subjects, i.e., plane geometry (1,746), solid geometry (332), and functions (534), where the latter two are all composed of newly collected problems. The choice of these three subjects is not only due to their rigorous demands on multi-modal reasoning, but also for two other considerations. For one thing, as we specialize MathVerse in mathematical problem-solving, other peripheral tasks in MathVista Lu2023MathVistaEM are not included, e.g., statistical reasoning, table question-answering, and puzzle tests. For another, we expect the evaluation to fully display the reasoning capabilities of MLLMs with moderate-level mathematical knowledge. This avoids limiting their performance with overly complex domain-specific theorems or prior commonsense knowledge. Therefore, we deliberately focus the collected problems on the high school level, excluding advanced college-level disciplines like calculus and graph theory featured in MMMU yue2023mmmu . Furthermore, expert annotators subdivide the problems into twelve fine-grained categories, as depicted in Figure 4, showcasing various dimensions of visual mathematical skills.
@@
[…]
@@ 2.2 Whether MLLMs Truly See the Diagrams?
Figure 6: Examples of the CoT Evaluation Strategy for MathVerse. We present two outputs from Qwen-VL-Max bai2023qwen with our CoT evaluation strategy, which assesses the fine-grained reasoning capabilities with a detailed explanation for error analysis.
@@ 2.3 CoT Evaluation Strategy
Compared to visual question-answering in general scenarios, the solving process of MLLMs for mathematical problems requires nuanced, step-by-step CoT reasoning. Considering two cases in Figure 6, one arrives at the correct solution albeit through incorrect intermediary steps, while the other demonstrates the opposite phenomenon. Therefore, the binary ‘Correct’ or ‘Incorrect’ evaluative approach of existing benchmarks is inadequate to examine the depth and precision of the multi-step reasoning process. To this end, we propose a CoT evaluation strategy to thoroughly assess their mathematical CoT skills in visual contexts, involving two prompting phases with GPT-4(V) openai2023gpt4v ; OpenAI2023GPT4TR .
@@ 3 Experiments
Table 2: Mathematical Evaluation on Six Problem Versions in MathVerse’s testmini Set. We calculate the ‘All’ score without averaging the ‘Text Only’ version. ‘CoT-E’ or ‘w/o’ denotes whether to employ the proposed CoT evaluation strategy or not. The highest accuracy for closed-source and open-source MLLMs is marked in red and blue respectively.
Model All Text Dominant Text Lite Text Only  Vision Intensive  Vision Dominant  Vision Only CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o CoT-E w/o Baselines Random Chance - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 Human - 64.9 - 71.2 - 70.9 - 41.7 - 61.4 - 68.3 - 66.7 LLMs ChatGPT ouyang2022training - - 51.3 33.3 38.5 18.9 51.3 33.3 - - - - - - GPT-4 OpenAI2023GPT4TR - - 63.4 46.5 40.7 20.7 63.4 46.5 - - - - - - Closed-source MLLMs Qwen-VL-Plus bai2023qwen 21.3 11.8 26.0 15.7 21.2 11.1 25.2 14.5 18.5 9.0 19.1 13.0 21.8 10.0 Gemini-Pro team2023gemini 35.3 22.3 39.8 27.6 34.7 23.7 44.5 27.9 32.0 19.4 36.8 20.3 33.3 20.5 Qwen-VL-Max bai2023qwen 37.2 24.8 42.8 30.3 37.7 24.8 47.9 32.2 33.6 20.6 35.9 23.3 35.9 25.1 GPT-4V openai2023gpt4v 54.4 38.3 63.1 52.1 56.6 40.9 60.3 46.1 51.4 34.9 50.8 33.6 50.3 29.8 Open-source MLLMs LLaMA-Adapter V2 gao2023llamaadapterv2 5.8 5.7 7.8 6.2 6.3 5.9 3.9 2.7 6.2 6.1 4.5 4.2 4.4 6.1 ImageBind-LLM han2023imagebind 10.0 9.3 13.2 11.4 11.6 11.3 12.9 11.7 9.8 8.9 11.8 11.2 3.5 3.4 mPLUG-Owl2 ye2023mplugowl2 10.3 4.6 11.6 6.6 11.4 6.3 13.8 6.1 11.1 6.3 9.4 5.6 8.0 4.9 MiniGPT-v2 chen2023minigpt 10.9 11.0 13.2 12.1 12.7 12.0 15.3 11.7 11.1 13.1 11.3 10.3 6.4 7.4 LLaVA-1.5 liu2023improvedllava 12.7 7.6 17.1 8.8 12.0 7.6 22.6 11.5 12.6 7.4 12.7 7.4 9.0 6.9 SPHINX-Plus gao2024sphinx 14.0 12.2 16.3 13.9 12.8 11.6 15.8 14.9 12.9 11.6 14.7 13.5 13.2 10.4 G-LLaVA gao2023g 15.7 16.6 22.2 20.9 20.4 20.7 21.6 21.1 16.5 17.2 12.7 14.6 6.6 9.4 LLaVA-NeXT liu2024llavanext 17.2 10.3 21.6 12.8 19.7 12.0 25.1 9.9 17.6 10.7 14.9 9.7 12.1 6.3 ShareGPT4V Chen2023ShareGPT4VIL 17.4 13.1 21.8 16.2 20.6 16.2 14.6 6.6 18.6 15.5 16.2 13.8 9.7 3.7 SPHINX-MoE gao2024sphinx 22.8 16.8 33.3 26.2 21.9 17.4 40.7 26.7 21.1 16.7 19.6 12.5 18.3 11.1 InternLM-XC2. dong2024internlm 25.9 16.3 36.9 20.2 28.3 14.3 42.5 24.5 20.1 14.2 24.4 17.5 19.8 15.2
@@ 3.1 Experimental Setup
We examine the performance of foundation models across three distinct categories on MathVerse: (a) Large Language Models (LLMs) as the text-only baseline, which only take textual questions as input, including ChatGPT OpenAI2023ChatGPT and GPT-4 OpenAI2023GPT4TR , (b) Closed-source MLLMs, represented by models like GPT-4V openai2023gpt4v , Gemini-Pro team2023gemini , Qwen-VL-Max bai2023qwen , and Qwen-VL-Plus, and (c) Open-source MLLMs, featuring models such as LLaVA-1.5 liu2023improvedllava (Vicuna-13B vicuna2023 ), LLaVA-NeXT liu2024llavanext (Vicuna-13B), SPHINX-MoE gao2024sphinx (Mixtral-8×7B albert24mixtral ), SPHINX-Plus (LLaMA2-13B touvron2023llama2 ), InternLM-XComposer2 dong2024internlm (InternLM2-7B team2023internlm ), LLaMA-Adapter V2 gao2023llamaadapterv2 (LLaMA-7B touvron2023llama ), ImageBind-LLM han2023imagebind (LLaMA-7B), MiniGPT-v2 chen2023minigpt (LLaMA2-7B), mPLUG-Owl2 ye2023mplugowl2 (LLaMA-7B), G-LLaVA gao2023g (LLaMA2-7B), and ShareGPT-4V Chen2023ShareGPT4VIL (Vicuna-13B).
@@
Model All Plane Geometry Solid Geometry Functions All Len Area Angle Anal Apply All Len Area Vol All Coord Prop Exp Apply Closed-source MLLMs Qwen-VL-Plus bai2023qwen 21.3 17.3 19.1 16.4 16.1 23.6 13.2 24.8 18.1 18.7 33.4 31.3 52.5 25.1 10.8 50.3 Gemini-Pro team2023gemini 35.3 33.0 32.2 42.6 28.4 30.2 32.3 33.4 35.0 29.3 36.1 28.3 25.7 26.6 10.8 51.3 Qwen-VL-Max bai2023qwen 37.2 38.4 41.7 46.4 32.6 40.6 38.7 33.7 25.4 28.3 42.6 38.4 43.7 35.5 13.6 61.0 GPT-4V openai2023gpt4v 54.4 56.9 60.8 63.4 52.6 48.5 60.9 50.2 54.8 39.9 56.8 52.8 72.3 47.1 30.9 70.1 Open-source MLLMs LLaMA-Adapter V2  gao2023llamaadapterv2 5.8 5.9 4.0 5.9 6.6 13.4 3.3 4.6 5.3 3.1 5.7 6.2 6.7 6.1 4.5 7.9 ImageBind-LLM han2023imagebind 10.0 9.7 12.1 9.9 9.2 10.2 4.8 4.6 4.9 3.5 5.3 14.9 12.3 13.8 4.6 25.9 mPLUG-Owl2 ye2023mplugowl2 10.3 7.7 8.2 6.0 5.7 12.4 10.6 11.0 9.2 6.7 15.7 17.4 22.8 18.6 5.3 22.2 MiniGPT-v2 chen2023minigpt 10.9 11.6 10.0 9.8 14.3 9.1 11.8 1.7 2.2 1.6 0.5 11.2 4.2 15.7 4.0 21.1 LLaVA-1.5 liu2023improvedllava 12.7 11.8 13.1 15.1 9.7 9.4 13.2 10.6 12.1 8.7 11.6 14.8 18.8 12.7 9.5 23.7 SPHINX-Plus gao2024sphinx 14.0 14.4 14.2 10.5 14.1 16.5 16.8 7.0 7.2 6.1 7.6 17.9 11.1 19.1 6.3 27.7 G-LLaVA gao2023g 15.7 20.2 17.3 13.6 26.5 5.9 23.1 5.0 10.3 4.4 3.1 9.2 9.1 9.1 1.3 15.5 LLaVA-NeXT liu2024llavanext 17.2 15.9 14.8 13.1 16.3 17.7 17.8 19.6 33.3 11.7 12.6 23.1 24.5 23.4 8.0 33.1 ShareGPT4V Chen2023ShareGPT4VIL 17.4 16.9 16.2 17.9 16.9 12.2 21.1 15.0 13.6 10.9 19.7 20.2 19.9 22.2 8.4 25.8 SPHINX-MoE gao2024sphinx 22.8 24.5 26.3 28.4 21.1 26.6 24.4 15.8 9.4 10.7 26.3 19.5 23.5 19.3 9.2 30.3 InternLM-XC2. dong2024internlm 25.9 26.2 27.1 29.7 20.6 18.5 22.2 20.1 34.5 14.1 25.2 23.7 24.4 24.9 10.6 36.3
@@
All our experiments are conducted under a zero-shot setting, showcasing the generalization capacity of MLLMs for mathematical reasoning, without few-shot prompting or further fine-tuning. By default, we employ the Chain-of-Thought (CoT) prompting technique wei2022chain , which encourages MLLMs to perform complete reasoning steps for a fine-grained evaluation. A baseline representing random chance is established for comparison, for which we select one option at random for multiple-choice questions and utilize empty for free-form questions. In addition, we recruit ten qualified college students, and ask them to solve the problems in MathVerse independently, serving as a baseline for human performance. We conduct all experiments on NVIDIA A100 GPUs. As the text-only LLMs can only take text questions as input, we evaluate them with the first three problem versions, i.e., Text Dominant, Text Lite, and Text Only. For the ‘w/o’ results, we utilize the template in MathVista Lu2023MathVistaEM to prompt GPT-4 OpenAI2023GPT4TR for answer extraction, and directly score the final answer without the intermediate reasoning process.

### v2

@@ 1 Introduction
With the substantial advances of big data and computational power, Large Language Models (LLMs) [4, 28, 55, 56, 13], such as ChatGPT [45] and GPT-4 [46], have emerged as a central point of interest in both industry and academia. To broaden their applicability across diverse contexts, Multi-modal Large Language Models (MLLMs) [66, 20, 52, 11, 61, 70] have recently become a fast-evolving track, exemplified by the latest GPT-4V [47], Gemini [22], and the open-source LLaVA [39, 34, 32, 30] and SPHINX [36, 21]. Concurrently, a diverse array of evaluation benchmarks [17, 40, 33, 18, 53] are curated to assess their visual comprehension performance across different domains. Notably, the capability to solve mathematical problems involving diagrams serves as a critical measure, offering insights into the multi-modal logical thinking prowess of MLLMs. This task demands MLLMs to accurately decode the visual elements within input diagrams (characters and figures), and correlate them with the condition specified by textual questions for mathematical reasoning. Previous efforts [42, 51], e.g., GeoQA [9, 5] and UniGeo [7], concentrate on the challenging geometric problems, while the recent MathVista [41] and MMMU [63] expand the scope to encompass broader disciplines, including functions, charts, and scientific problems.
Figure 1: (a) We showcase three examples of Text Redundancy (highlighted in red) within existing visual math benchmarks [9, 41, 63]. (b) We report an ablation study by respectively removing the redundant texts and input diagrams on 120 randomly selected problems, for closed-sourced [47, 22, 3] and open-sourced [21, 38, 16] MLLMs.
@@
Figure 2: Comparison of Visual Mathematical Reasoning by Three MLLMs. Despite the incorrect final answer, GPT-4V [47], Gemini-Pro [22], and SPHINX-MoE [21] exhibit different levels of quality in the intermediate reasoning process.
@@
Second, to rigorously assess the visual Chain-of-Thought (CoT) capabilities [58], we propose a CoT Evaluation strategy for the step-by-step reasoning assessment of MLLMs. For each model’s output, we leverage GPT-4 to first extract several crucial steps exclusively from the solving process, deliberately omitting the input of the question and answer. This approach aims to mitigate the bias towards GPT-4’s inherent question-answering propensities. Then, the corresponding question, diagram, and ground-truth answer are fed into GPT-4 to evaluate each identified critical step, and provide detailed error analysis. Finally, the overall score is obtained by considering every single step within reasoning. Note that, we do not pre-define a ground-truth key-step template, since each math problem may encompass a variety of solution pathways, and different MLLMs tend to exhibit variable reasoning lengths. With CoT scoring, MathVerse showcases a fine-grained evaluation of the intermediate logical deduction of MLLMs, demonstrating visual mathematical CoT capabilities.
@@
We conduct extensive experiments on MathVerse with popular closed-source [47, 3, 22] and open-source [37, 38, 16, 21] MLLMs. Comparing different problem versions, we unveil that, most existing MLLMs struggle to understand math diagrams, relying heavily on textual questions. Therein, GPT-4V [47] achieves the best overall performance across different problem versions and subjects. Surprisingly, some of the MLLMs even attain much higher results without the diagram input, e.g., +5.1% for Qwen-VL-Max [3] and +5.6% for InternLM-XComposer2 [16]. With the fine-grained error analysis produced by our CoT evaluation strategy, we demonstrate such results are due to their deficient visual encoding capacity for mathematical diagrams, which instead acts as a distraction for problem-solving. In contrast, GPT-4V and ShareGPT4V [12] demonstrate relatively better comprehension of the visual content for mathematical reasoning. Our experimental results suggest that inadequate mathematical visual interpretation capabilities represent the most significant impediment for MLLMs in addressing multi-modal math problems, indicating substantial potential for advancement.
@@ 2 MathVerse
1https://homework.study.com2https://www.ixl.com/math3https://mathspace.co/us
@@ 2.1 Visual Math Dataset
MathVerse comprises a total of 2,612 visual math problems, which contribute to the final created 15K test samples. Detailed statistics for data composition are presented in Table 4. This meticulously collected dataset covers three fundamental math subjects, i.e., plane geometry (1,746), solid geometry (332), and functions (534), where the latter two are all composed of newly collected problems. The choice of these three subjects is not only due to their rigorous demands on multi-modal reasoning, but also for two other considerations. For one thing, as we specialize MathVerse in mathematical problem-solving, other peripheral tasks in MathVista [41] are not included, e.g., statistical reasoning, table question-answering, and puzzle tests. For another, we expect the evaluation to fully display the reasoning capabilities of MLLMs with moderate-level mathematical knowledge. This avoids limiting their performance with overly complex domain-specific theorems or prior commonsense knowledge. Therefore, we deliberately focus the collected problems on the high school level, excluding advanced college-level disciplines like calculus and graph theory featured in MMMU [63]. Furthermore, expert annotators subdivide the problems into twelve fine-grained categories, as depicted in Figure 4, showcasing various dimensions of visual mathematical skills.
@@
Our collection procedure for high-quality visual math problems involves a rigorous selection from both pre-existing datasets and public question repositories. In the domain of plane geometry, we initially select 750 problems from GeoQA [9], 119 from GEOS [51], and 507 from Geometry3K [42], based on their original data quality and distribution. We exclude questions that are extremely simple or excessively complex, as well as those that appear dubious or lack necessary conditions. To enhance the diversity of question types and diagram styles, we further enrich our dataset with additional 370 plane geometry problems by manually collecting from other sources1,2,3. Given the scarcity of solid geometry and function-related problems in existing benchmarks, we purposefully gather these two types of problems (332 and 534, respectively) from new sources1,2,3 to address this gap. Problems that include multiple diagrams or require visual illustrations within solutions are excluded, considering the current limitations of MLLMs in resolving such information. Note that, all the newly collected problems (1,236) accompany detailed explanations. After the preliminary collection, we undertake a comprehensive review to verify the accuracy of the answers, ensure consistency between questions and diagrams, and confirm the relevance of each problem to the defined twelve categories. This meticulous review guarantees the dataset’s quality and precision.
@@ 2.2 Whether MLLMs Truly See the Diagrams?
Figure 6: Examples of the CoT Evaluation Strategy for MathVerse. We present two outputs from Qwen-VL-Max [3] with our CoT evaluation strategy, which assesses the fine-grained reasoning capabilities with a detailed explanation for error analysis.
@@ 2.3 CoT Evaluation Strategy
Compared to visual question-answering in general scenarios, the solving process of MLLMs for mathematical problems requires nuanced, step-by-step CoT reasoning. Considering two cases in Figure 6, one arrives at the correct solution albeit through incorrect intermediary steps, while the other demonstrates the opposite phenomenon. Therefore, the binary ‘Correct’ or ‘Incorrect’ evaluative approach of existing benchmarks is inadequate to examine the depth and precision of the multi-step reasoning process. To this end, we propose a CoT evaluation strategy to thoroughly assess their mathematical CoT skills in visual contexts, involving two prompting phases with GPT-4(V) [47, 46].
@@ 3 Experiments
Table 2: Mathematical Evaluation on Six Problem Versions in MathVerse’s testmini Set. We calculate the ‘All’ score without averaging the ‘Text Only’ version. ‘CoT-E’ or ‘Acc’ denotes whether to employ the proposed CoT evaluation strategy or not. The highest accuracy for closed-source and open-source MLLMs is marked in red and blue respectively.
Model All Text Dominant Text Lite Text Only  Vision Intensive  Vision Dominant  Vision Only CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc CoT-E Acc Baselines Random Chance - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 - 12.4 Human - 64.9 - 71.2 - 70.9 - 41.7 - 61.4 - 68.3 - 66.7 LLMs ChatGPT [48] - - 51.3 33.3 38.5 18.9 51.3 33.3 - - - - - - GPT-4 [46] - - 63.4 46.5 40.7 20.7 63.4 46.5 - - - - - - Closed-source MLLMs Qwen-VL-Plus [3] 21.3 11.8 26.0 15.7 21.2 11.1 25.2 14.5 18.5 9.0 19.1 13.0 21.8 10.0 Gemini-Pro [22] 35.3 23.5 39.8 26.3 34.7 23.5 44.5 27.3 32.0 23.0 36.8 22.3 33.3 22.2 Qwen-VL-Max [3] 37.2 25.3 42.8 30.7 37.7 26.1 47.9 28.9 33.6 24.1 35.9 24.1 35.9 21.4 GPT-4V [47] 54.4 39.4 63.1 54.7 56.6 41.4 60.3 48.7 51.4 34.9 50.8 34.4 50.3 31.6 Open-source MLLMs LLaMA-Adapter V2 [20] 5.8 5.7 7.8 6.2 6.3 5.9 3.9 2.7 6.2 6.1 4.5 4.2 4.4 6.1 ImageBind-LLM [24] 10.0 9.2 13.2 11.4 11.6 11.3 12.9 11.7 9.8 8.9 11.8 11.2 3.5 3.4 mPLUG-Owl2 [62] 10.3 5.9 11.6 6.6 11.4 6.3 13.8 6.1 11.1 6.3 9.4 5.6 8.0 4.9 MiniGPT-v2 [11] 10.9 11.0 13.2 12.1 12.7 12.0 15.3 11.7 11.1 13.1 11.3 10.3 6.4 7.4 LLaVA-1.5 [37] 12.7 7.6 17.1 8.8 12.0 7.6 22.6 11.5 12.6 7.4 12.7 7.4 9.0 6.9 SPHINX-Plus [21] 14.0 12.2 16.3 13.9 12.8 11.6 15.8 14.9 12.9 11.6 14.7 13.5 13.2 10.4 G-LLaVA [19] 15.7 16.6 22.2 20.9 20.4 20.7 21.6 21.1 16.5 17.2 12.7 14.6 6.6 9.4 LLaVA-NeXT [38] 17.2 15.6 21.6 19.4 19.7 15.2 25.1 18.1 17.6 16.8 14.9 15.2 12.1 11.3 ShareGPT4V [12] 17.4 13.1 21.8 16.2 20.6 16.2 14.6 6.6 18.6 15.5 16.2 13.8 9.7 3.7 SPHINX-MoE [21] 22.8 15.0 33.3 22.2 21.9 16.4 40.7 18.3 21.1 14.8 19.6 12.6 18.3 9.1 InternLM-XC2. [16] 25.9 16.5 36.9 22.3 28.3 17.0 42.5 16.5 20.1 15.7 24.4 16.4 19.8 11.0
@@ 3.1 Experimental Setup
We examine the performance of foundation models across three distinct categories on MathVerse: (a) Large Language Models (LLMs) as the text-only baseline, which only take textual questions as input, including ChatGPT [45] and GPT-4 [46], (b) Closed-source MLLMs, represented by models like GPT-4V [47], Gemini-Pro [22], Qwen-VL-Max [3], and Qwen-VL-Plus, and (c) Open-source MLLMs, featuring models such as LLaVA-1.5 [37] (Vicuna-13B [13]), LLaVA-NeXT [38] (Vicuna-13B), SPHINX-MoE [21] (Mixtral-8×7B [28]), SPHINX-Plus (LLaMA2-13B [56]), InternLM-XComposer2 [16] (InternLM2-7B [54]), LLaMA-Adapter V2 [20] (LLaMA-7B [55]), ImageBind-LLM [24] (LLaMA-7B), MiniGPT-v2 [11] (LLaMA2-7B), mPLUG-Owl2 [62] (LLaMA-7B), G-LLaVA [19] (LLaMA2-7B), and ShareGPT-4V [12] (Vicuna-13B).
@@
Model All Plane Geometry Solid Geometry Functions All Len Area Angle Anal Apply All Len Area Vol All Coord Prop Exp Apply Closed-source MLLMs Qwen-VL-Plus [3] 21.3 17.3 19.1 16.4 16.1 23.6 13.2 24.8 18.1 18.7 33.4 31.3 52.5 25.1 10.8 50.3 Gemini-Pro [22] 35.3 33.0 32.2 42.6 28.4 30.2 32.3 33.4 35.0 29.3 36.1 28.3 25.7 26.6 10.8 51.3 Qwen-VL-Max [3] 37.2 38.4 41.7 46.4 32.6 40.6 38.7 33.7 25.4 28.3 42.6 38.4 43.7 35.5 13.6 61.0 GPT-4V [47] 54.4 56.9 60.8 63.4 52.6 48.5 60.9 50.2 54.8 39.9 56.8 52.8 72.3 47.1 30.9 70.1 Open-source MLLMs LLaMA-Adapter V2  [20] 5.8 5.9 4.0 5.9 6.6 13.4 3.3 4.6 5.3 3.1 5.7 6.2 6.7 6.1 4.5 7.9 ImageBind-LLM [24] 10.0 9.7 12.1 9.9 9.2 10.2 4.8 4.6 4.9 3.5 5.3 14.9 12.3 13.8 4.6 25.9 mPLUG-Owl2 [62] 10.3 7.7 8.2 6.0 5.7 12.4 10.6 11.0 9.2 6.7 15.7 17.4 22.8 18.6 5.3 22.2 MiniGPT-v2 [11] 10.9 11.6 10.0 9.8 14.3 9.1 11.8 1.7 2.2 1.6 0.5 11.2 4.2 15.7 4.0 21.1 LLaVA-1.5 [37] 12.7 11.8 13.1 15.1 9.7 9.4 13.2 10.6 12.1 8.7 11.6 14.8 18.8 12.7 9.5 23.7 SPHINX-Plus [21] 14.0 14.4 14.2 10.5 14.1 16.5 16.8 7.0 7.2 6.1 7.6 17.9 11.1 19.1 6.3 27.7 G-LLaVA [19] 15.7 20.2 17.3 13.6 26.5 5.9 23.1 5.0 10.3 4.4 3.1 9.2 9.1 9.1 1.3 15.5 LLaVA-NeXT [38] 17.2 15.9 14.8 13.1 16.3 17.7 17.8 19.6 33.3 11.7 12.6 23.1 24.5 23.4 8.0 33.1 ShareGPT4V [12] 17.4 16.9 16.2 17.9 16.9 12.2 21.1 15.0 13.6 10.9 19.7 20.2 19.9 22.2 8.4 25.8 SPHINX-MoE [21] 22.8 24.5 26.3 28.4 21.1 26.6 24.4 15.8 9.4 10.7 26.3 19.5 23.5 19.3 9.2 30.3 InternLM-XC2. [16] 25.9 26.2 27.1 29.7 20.6 18.5 22.2 20.1 34.5 14.1 25.2 23.7 24.4 24.9 10.6 36.3
@@
All our experiments are conducted under a zero-shot setting, showcasing the generalization capacity of MLLMs for mathematical reasoning, without few-shot prompting or further fine-tuning. By default, we employ the Chain-of-Thought (CoT) prompting technique [58], which encourages MLLMs to perform complete reasoning steps for a fine-grained evaluation. A baseline representing random chance is established for comparison, for which we select one option at random for multiple-choice questions and utilize empty for free-form questions. In addition, we recruit ten qualified college students, and ask them to solve the problems in MathVerse independently, serving as a baseline for human performance. We conduct all experiments on NVIDIA A100 GPUs. As the text-only LLMs can only take text questions as input, we evaluate them with the first three problem versions, i.e., Text Dominant, Text Lite, and Text Only. For the ‘w/o’ results, we utilize the template in MathVista [41] to prompt GPT-4 [46] for answer extraction, and directly score the final answer without the intermediate reasoning process.